OpenStudy (anonymous):

Using the formal definition of limits, prove that lim(x -> -1.5) (9-4x^2)/(2x+3) = 6. So far I have: |((-2x+3)(2x+3))/(2x+3)-6| < e |-2x -3| < e But I don't know if I should do: |x+3/2|> -e/2 or to just leave it. Someone please help!

OpenStudy (jamesj):

That last step is wrong, because the negative never makes it outside the absolute value sign. Write f(x) = (9-4x^2)/(2x+3). Then what you have have is | f(x) - 6 | = | (-2x+3)(2x+3)/(2x+3) - 6 | = | -2x + 3 - 6 | = | 2x + 3 | = 2 | x - (-3/2) | Now let epsilon e be an arbitrary number e > 0. Then choose delta d = e/2. Then 0 < | x - (-3/2) | < d = e/2 ==> | f(x) - 6 | < e Therefore by definition of the limit: \[\lim_{x \rightarrow -3/2} \ \ f(x) = 6.\]

OpenStudy (jamesj):

Epsilon-delta proofs are probably the most subtle mathematics you've learnt so far. So do yourself a favor (and I'll feel better about it too!). Take a blank piece of paper and write out the proof above and make sure you understand every step, particularly at the end. When you can do that with comfort without looking at what we're written here, then you've really learnt something.

OpenStudy (anonymous):

Thanks James! I appreciate your help. I actually ended up figuring it out. I tend to make minor mistakes with my math, especially when I am tired. I understand it now though! :D

OpenStudy (anonymous):

Hey James, can you help me again? I don't know if I did the following question right: "Prove that \[\lim_{x \rightarrow -\infty} (1 + x^3) = -\infty\] using the precise definition of infinite limits (i.e. x < N then f(x) < M)." I did: 1 + x^3 < M x < (M -1)^(1/3) N > 0, so N = max(1, (M - 1)^(1/3)) x < N = max(1, (M - 1)^(1/3)) 1 + x^3 < max(1+x^3,M) < M Therefore, \[\lim_{x \rightarrow -\infty} (1 + x^3) = -\infty\]

OpenStudy (jamesj):

It is not the case that N > 0 and "1 + x^3 < max(1+x^3,M) < M" this step isn't quite right. Go through it again carefully.

OpenStudy (anonymous):

I am confused :|

OpenStudy (jamesj):

Go back to the definition. The idea here is that for all M < 0, we can find an N such that \[x < N \implies f(x) < M\] So where exactly is the clean logic here. Show me how to start with an M and find such an N that gives us f(x) < M. You've almost got it, but you have a couple of non sequitors in your logic, such as the two I wrote above.

OpenStudy (anonymous):

Oh, I thought M and N were supposed to be positive, but they are supposed to be negative?

OpenStudy (jamesj):

clearly because we're taking the limit as x --> NEGATIVE infinity and the limit we want is also negative infinity.

OpenStudy (jamesj):

So we want to get f(x) arbitrarily large and negative provided x is sufficiently large and negative.

OpenStudy (anonymous):

Okay, so do I have up to x < (M-1)^(1/3)) correct?

OpenStudy (jamesj):

Yes, that's all fine.

OpenStudy (jamesj):

and that's the preamble you want to knowing how to choose N given M

OpenStudy (anonymous):

Okay So then N < 0.. so do we still take max(1, (M-1)^(1/3))? I know I am lost here, but my teacher hasn't shown us how to do this and I watched a video on x approaching positive infinity, and I got confused.

OpenStudy (jamesj):

Hence ... given M < 0, choose N = (something involving M). Then x < N ==> f(x) < M and therefore by definition the limit as x --> -infty of f(x) = -infty.

OpenStudy (jamesj):

So given M < 0, what is your choice for N which will give us the logical deduction we want?

OpenStudy (anonymous):

I'm not sure what you are asking

OpenStudy (jamesj):

It's exactly like in the epsilon-delta proof. Given epsilon, what choice of delta will give us 0 < |x-a| < delta ==> | f(x) - L | < epsilon

OpenStudy (anonymous):

Okay... Could you walk me through this? We aren't going to be taught in class because we are behind, and I am obviously not getting it.

OpenStudy (jamesj):

Here, the formulation of the limit is this: if for all M < 0 (i.e., large and negative) we can find an N such that x<N ==> f(x)<M Then we will say that is the definition that \[\lim_{x \rightarrow -\infty} \ f(x) = -\infty\]

OpenStudy (anonymous):

Okay, I get that much

OpenStudy (jamesj):

Ok, so given M , what choice of N should we make such that x<N ==> f(x)<M

OpenStudy (jamesj):

Given M < 0, let N = some expression involving M, then x<N ==> f(x)<M

OpenStudy (jamesj):

What is that expression for N?

OpenStudy (anonymous):

Is that what I solved for? As in (M - 1)^(1/3)?

OpenStudy (jamesj):

Yes. Given M < 0, let N = (M-1)^(1/3), then x < N ==> x < (M-1)^(1/3) ==> 1 + x^3 < M Hence by definition ....

OpenStudy (anonymous):

Oh! That makes sense So, by definition the limit is -infinity... if that is what you were asking?

OpenStudy (jamesj):

No, that's what YOU are asking. And yes.

OpenStudy (jamesj):

What I don't blame you for getting confused about, is bounding these numbers to be large, negative.

OpenStudy (anonymous):

Thanks so much! I know I can be a bother. I just feel lost at times when my teacher jumps over things. You are a great help!

OpenStudy (jamesj):

But work this example again and make up your own example for the other three types of limits x goes to plus/minus infinity and the limit is plus/minus infinity.

OpenStudy (jamesj):

Figure out exactly what are the definitions, and then make up some of your own examples and work them through.

OpenStudy (anonymous):

I will :)

OpenStudy (jamesj):

Now, I think I deserve a medal.

OpenStudy (anonymous):

Of course I forget about those

OpenStudy (anonymous):

James, could you help me with the following question? Please! Show that det(A) =1/2 |tr(A) 1| |tr(A^2) tr(A)| for every 2 x 2 matrix.

OpenStudy (jamesj):

Doing some work of my own right now ... give me a few

OpenStudy (anonymous):

No problem :)

OpenStudy (jamesj):

OK, so writing A = a b c d then A^2 = a^2 + bc ab + bd ac + cd bc + d^2 Hence tr(A^2) = a^2 + 2bc + d^2

OpenStudy (jamesj):

and tr(A) = a + d of course Now the determinant of that matrix is tr(A)^2 - tr(A^2) = (a+d)^2 - (a^2 + 2bc + d^2) = 2 (ad - bc) = 2 det(A)

OpenStudy (jamesj):

So this is straightforward, certainly much easier than your limit problems.

OpenStudy (jamesj):

And frankly, I'd bet a dollar this is a completely useless identity. I think the question was just made up to get your comfortable manipulating these expressions from matrices

OpenStudy (anonymous):

I believe so :) Thank you very much! You are a great help :)

OpenStudy (anonymous):

Hey James. I have another quick problem that I am a little confused about. Our prof wanted us to attempt it before our lesson, but that is not going well. I was wondering if you could help me with it? Consider the vector v = (1,-5) a) Find all values of k such that u = (k,7) is orthogonal to v b) Find all values of k such that u = (k,7) is parallel to v For a, I computed the dot product because I know that when two vectors are perpendicular they have a dot product of zero. So I found that when k = 35, u is orthogonal to v. I am not sure if that finds all values of k, or if I chose the right method. But I think it works. I am not sure how to attempt b. I know that two parallel vectors are scalar multiples of each other, but I am little confused.

OpenStudy (jamesj):

So for a, you're exactly right: the two vectors are orthogonal if and only if (k,7).(1,-5) = 0 <=> k - 35 = 0 <=> k = 35 and hence k = 35 is the only value of k for which they are orrthongal.

OpenStudy (jamesj):

Now (k,7) is parallel to (1,-5) if there exists a nonzero scalar, c such that (k,7) = c(1,-5) <=> (k,7) = (c,-5c) Now that is the case if both k = c and 7 = -5c By the second of those equation c = -7/5 and now by the first k = -7/5 also.

OpenStudy (anonymous):

Oh, so u is parallel to v if k = -7/5? That makes sense. I remember doing something like that in grade 12. Thanks for your help!

OpenStudy (jamesj):

sure

OpenStudy (anonymous):

Hey James, when you have a moment, would you mind looking at my other question? Please and thank you!